OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 5x – 3y = 1 7x – 4y = 2

OpenStudy (anonymous):

1. Elimination method is where you manipulate the equations so that there is an equal number of x's in both equations or an equal number of y's. Then you can add or subtract the equations to eliminate that variable. Part 2: 20x - 12y = 4 21x - 12y = 6 therefore x = 6-4 = 2 So x = 2, and y = 3.